For a body moving in circular motion with a constant angular velocity, the magnitude of average acceleration over a period of half a revolution is x times th emagnitude of its instantaneous acceleration. Find x.

Grade:11

1 Answers

Vikas TU
7 years ago
Average accln. = (vjcap – (-vjcap))/(pi/w) = > 2vjcap*w/pi
where w is constant angular velocity given.
Now, Given
2v*w/pi  =  x * dv/dt
dv/v  = (2*w/x*pi)dt
integerate both sides
logv = 2wt/pi*x
x = 2wt/(pi)logv